OpenStudy (anonymous):

Write a proportion for the following problem: Roberta jogs 2.2 miles in 25 minutes. At that rate, how many miles could she jog in 33 minutes?

OpenStudy (anonymous):

2.2/25=x/33

OpenStudy (anonymous):

x=33*2.2/25=2.904

OpenStudy (anonymous):

give me a medal

OpenStudy (anonymous):

2.2/x = 33/25 x/2.2 = 25/33 25/2.2 = x/33 2.2/25 = x/33